OpenStudy (yanasidlinskiy):

Can someone help me? Simplify the trigonometric expression. (Give me a sec)

OpenStudy (yanasidlinskiy):

\[\frac{ 1 }{ 1+\sin \Theta }+\frac{ 1 }{ 1-\sin \Theta } \]

OpenStudy (yanasidlinskiy):

Would it be \[2\cos ^{2} \Theta\] ??

OpenStudy (anonymous):

i think the answer would be 2/cos^2 theta

OpenStudy (yanasidlinskiy):

That isn't in ANY of my answer choices.

OpenStudy (anonymous):

is 2sec^2 theta one of your answer choices?

OpenStudy (yanasidlinskiy):

There is NO fractions:)

OpenStudy (yanasidlinskiy):

Yes. So..I'm wrong?

OpenStudy (anonymous):

so i think the answer is 2sec^2 x

OpenStudy (yanasidlinskiy):

Thanx:) I'll have to double check that though:) I'm not saying that you're wrong though. I just don't see why it's "sec" not "cos"

OpenStudy (yanasidlinskiy):

@ikram002p

OpenStudy (anonymous):

um sec^2 x = 1/cos^2 x. its a standard trig identity :)